Round Percent Change to the nearest tenth of a percent
ORIGINAL AMOUNT: CHANGE: PERCENT CHANGE: NEW AMOUNT:
1.)__________ ________ 150% increase 48
2.)__________ add 0.3 125% increase ________

Answer by Theo(13342)

Round Percent Change to the nearest tenth of a percent
ORIGINAL AMOUNT: CHANGE: PERCENT CHANGE: NEW AMOUNT:

1.)__________ ________ 150% increase 48

150% increase means you increased the original amount by 150% of the original amount.

x = original amount
y = new amount
150% of the original amount means 1.5 * the original amount.

formula would be:
y = x + 1.5 * x
if y = 48, then:
48 = x + 1.5 * x
combine like terms to get:
48 = 2.5 * x
divide both sides of this euation by 2.5 to get:
x = 48 / 2.5 = 19.2

original amount is 19.2.
you increase it by 150% of 19.2 = 1.5 * 19.2 = 28.8
19.2 + 28.8 = 48




ORIGINAL AMOUNT: CHANGE: PERCENT CHANGE: NEW AMOUNT:

2.)__________ add 0.3 125% increase ________

.3 increase means you increased the original amount by 0.3,
.3 represents a 125% increase over the original amount.
x = original amount
y = new amount
.3 = 125% of the original amount means that .3 = 1.25 * the original amount.

formula would be:
.3 = 1.25 * x
divide both sides of this equation by 1.25 to get:
.3 / 1.25 = x
solve for x to get:
x = .3 / 1.25 = .24

increase of .3 is equal to 125% of original amount of .24 = 1.25 * .24 = .3

formula for new amount = y would be:
y = x + .3
since x = .24, then:
y = .24 + .3 which results in:
y = .54

original amount is .24
you increase it by 125% of .24 = 1.25 * .24 = .3
new amount is equal to .24 + .3 = .54